OpenStudy (anonymous):

The scale factor of a model of a hot air balloon to the actual hot air balloon is 1 to 5. The volume of the actual balloon is 5000 m3. What is the volume of the model?

OpenStudy (anonymous):

1000

Directrix (directrix):

Solve this for x: (1/5)^3 = x/5000

Directrix (directrix):

1/125 = x/5000 Cross Multiply 125x = 5000 Divide both sides by 125 x = ? @katrinamarie0803

OpenStudy (anonymous):

40?

Directrix (directrix):

Yes, that is what I got. 40 m3. @katrinamarie0803